OpenStudy (anonymous):

Suppose n is any positive integer having an even number of digits, and m is its reverse, for example if n = 1234, then m = 4321. a) Show that m+n is always divisible by 11. b) Why doesn't the same proof work for positive integers having an odd number of digits?

OpenStudy (anonymous):

*bookmark* :D

OpenStudy (anonymous):

This is based on idea of 10a+b+10b+a = 11(a+b) for 2 digit numbers..I guess u can proceed inductively. You need pairs, so it won't work for odd (getting a counterexample for odd n is easy).

OpenStudy (anonymous):

Okay, but, for 4 digit numbers, you have 1000a+b+1000b+a = 1001(a+b), and for 6 digits, you have 100001(a+b)... and so on. So how do you know that these numbers are divisible by 11 as well?

OpenStudy (anonymous):

Not sure if I understand u, u mean how do I "know" 1001" is divisible by 11? That's why I said proceed inductively, for 4 digits, 1001a +110b +110c +1001d, etc. Looks like u want 10^(2n-1) +1 / 11 for all n, use induction.

OpenStudy (anonymous):

I know the answer, I invented the question myself... I just want to see if you can write the induction proof.

OpenStudy (anonymous):

Yes, I can write the induction proof, thats why I said @use induction@